5) Assume that A)00 + 5t. (1)-| a) Find i,. b) Find i 10.
Answer #1

i = A(t) - A(t-1) / A(t-1)

a) i5

=[100+5(5) - 100-5(4)] / 100 + 5(4)

=5/120

=0.04166

b)i10

=[100+5(10) - 100-5(9)] / 100 + 5(9)

=5/145

=0.03448